# VisLTR: Visualization-in-the-Loop Table Reasoning

### v1

@@ Abstract
Table reasoning transforms user requirements into corresponding answers according to the provided table, which is often integrated with natural language interfaces for lay users to explore tabular data effortlessly. Recent research exploits large language models to facilitate table reasoning, by transforming vague user requirements into structured query languages (SQLs). However, these SQL-based approaches often overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries. To this end, VisLTR is designed as a visualization-in-the-loop table reasoning framework that leverages visualizations as a proxy to provide concise data representations, capture interesting data patterns, and support cross-modal analysis. We describe VisLTR as a process consisting of four major modules: 1) visualization alignment that utilizes large vision-language models to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multi-modal interactions for user-friendly table reasoning. Quantitative evaluation demonstrates the effectiveness of the alignment model in cross-modal visualization pairings. We further demonstrate applications of the framework on various table reasoning tasks such as table summarization and pattern detection.
@@
Table reasoning, visualization reference, multi-modal learning
@@
VisLTR supports data change pattern identification and multi-modal explorations of text queries and hand-drawn sketches using visualizations as a proxy between a table and user requirements (bottom). The proposed approach advances existing SQL-based LLM methods for table reasoning that overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries (top).
@@ 1 Introduction
Table reasoning is the task of deriving answers from tables (i.e., tabular data) based on specific questions[58], supporting a wide range of applications such as fact verification[6] and question answering[21]. The key challenge of table reasoning lies in understanding both the intent of queries (e.g., a text query) and the semantics of tables, and accurately generating answers that require logical reasoning. Existing methods for table reasoning can be broadly categorized into generic table reasoning that directly generate answers using pre-trained models or large language models (LLMs); and program-aided table reasoning that utilizes executable query languages (e.g., SQL or SPARQL) as the bridge between user queries and answers derived from the table[49]. Generic table reasoning performs reasoning implicitly[49, 5], which, however, is poor at tracking how answers are derived and increases the risk of plausible but incorrect responses. Instead, program-aided table reasoning ensures precise data retrieval and offers transparency in the reasoning process, which is essential for applications such as fact verification[56, 7]. Therefore, program-aided reasoning methods are attracting significant attention from both academia and industry[15].
@@
Despite the growing popularity of program-aided reasoning methods, there exist several limitations. One of the primary issues stems from SQL’s inherent lack of analytical functionality[24], which leads to an overlook of data change patterns. Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A) illustrates a scenario where users seek to examine the price trend of Apple during March. Despite successfully retrieving a sub-table containing the price of Apple in March, SQL-based LLMs fall short when it comes to further querying and discerning change patterns within this sub-table. Second, LLMs are known to encounter drift issues, which refer to the loss of information in the middle of long contexts. LLM drift is common when processing large data tables[4], where the LLM’s reasoning process may encounter errors due to the loss of information in the middle, as illustrated in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B). Last, table reasoning is often integrated with natural language interfaces, to facilitate lay users’ exploration of tabular data effortlessly. However, there are scenarios in which laymen are unable to express their intentions in natural language but would prefer to use hand-drawn sketches. An example is shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), where a user may not know the term “two-valley” but can easily sketch this pattern. Existing SQL-based LLMs do not support non-verbal interactions, which are highly demanded.
@@
To overcome the limitations, we introduce a visualization-in-the-loop table reasoning framework that incorporates visualization references into large vision-language models (LVLMs) to complement SQL-based LLMs for table reasoning. Our key idea is to transform a tabular data into a set of meaningful and insightful visualizations, termed visualization references, which serve as a proxy to capture data insights and align with human intentions. This framework has several advantages compared to existing SQL-based table reasoning methods. First, visualization references can capture the short- and long-term data change patterns. Considering the price trend query as an example (Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A)), the retrieved visualization reference could capture the “two-peak” price trend of Apple during March. Second, regardless of the length of data facets, the encapsulated visualization references can capture and retain the data change patterns in fixed-size images. Adopting fixed-size visualization references as the representation of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning. As Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B) shows, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference. Moreover, leveraging visualizations as a proxy offers a new modality of hand-drawn sketches for interacting with the tabular data, which align more intuitively with human cognition[27]. When querying for a hand-drawn sketch as shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), it is easier to retrieve a similar visualization reference among the candidates than to retrieve a similar change pattern in the tabular data[35].
@@
However, the strategy presents several challenges. First, data tables often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references. (C1) Ensuring the visualization references are comprehensive enough to accurately represent an input data table, thereby meeting the diverse analytical needs, poses the first challenge. Second, (C2) effectively aligning textual descriptions, visualization images, and hand-drawn sketches to bridge user queries with visualization references, meanwhile enabling multi-modal interactions presents a challenge. Existing models for vision-language tasks such as chart question answering (CQA)[19] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns. Last, to comprehensively represent a table, a large number of visualization references will be generated. (C3) Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge.
@@
To evaluate the effectiveness of the proposed framework, we design a prototype system, which encompasses four main modules aimed at addressing the challenges C1--C3. The visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table. To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1). Second, in the visualization alignment module, we fine-tune a large vision-language model (LVLM) to connect visualizations of various modalities including charts, text, and sketches (addressing C2). To fine-tune the model, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions. We further leverage the CLIP model[41] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment. Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3). Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multi-modal interactions. Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisLTR with the input table. Specifically, a user query is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned LVLM. Subsequently, these descriptions are filled back into the input table to generate textual responses for the user. Quantitative evaluations demonstrate the effectiveness of the fine-tuned LVLM, and case studies illustrate the system’s applicability in diverse application scenarios. Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
@@
(1) Visualization-in-the-Loop Framework. We introduce a visualization-in-the-loop framework that leverages visualizations as the proxy to support table reasoning. The framework complements existing program-aided approaches for table reasoning, particularly in capturing data change patterns and facilitating non-verbal exploration.
(2) Multi-modal Visualization Alignment. We design an effective LVLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches. The model aligns multi-modal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
(3) VisLTR System. We develop a prototype system that supports the exploration of large-scale and multiple time-series tables through intuitive user interactions. Various case studies and quantitative results prove the effectiveness and applicability of VisLTR.
@@ 2.1 Table Reasoning
Table reasoning integrated with natural language (NL) interfaces supports lay users to explore tabular data effortlessly. This capability is essential for many complex tasks such as fact verification[6] and question answering[21]. Existing table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[49]. Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[5]. Due to the complexity of table structure, these methods face difficulties in dealing with large tables. Moreover, explainability is limited due to their end-to-end nature. To facilitate complex table reasoning, program-aided table reasoning utilizes executable query languages, such as SQL, to form a more structured and accountable bridge between user quires and the derived answers[11, 31, 20]. Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[56], using prompt engineering[25], chain-of-thought approaches[49, 57], and pre-training techniques[48]. However, SQL-based LLM table reasoning approaches encounter intricate limitations due to their reliance on SQL. First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[24], overlooking data change patterns. Second, when confronted with large tables, LLMs may encounter drift issues[4] while accessing information in the middle of long contexts. Moreover, in many scenarios, relying solely on textual inputs restricts the expression of non-verbal intentions, such as hand-drawn sketches.
@@
To overcome these limitations, this work introduces a novel framework that leverages visualizations as a proxy between tables and user queries, complementing SQL-based table reasoning by addressing challenges such as data change pattern identification and multi-modal interaction. The strategy is feasible, as visualizations can depict intuitive data patterns and illustrate complex underlying processes[14], facilitating swift and effective interpretation by the human brain[26]. However, this approach also presents challenges to be tackled. First, tabular data typically encompass numerous features, leading to a plethora of potential visualizations[54, 46] that complicate interactive interactions. Second, there are no direct approaches available to connect free-form user queries with visualizations.
Figure 1: An overview of the proposed framework, which mainly contains four main modules: 1) the visualization referencing module decomposes a raw data table into facets and maps them to visualization references; 2) the visualization pruning module leverages pruning and indexing techniques to ease the storage and improve the retrieve efficiency; 3) the visualization alignment module leverages a fine-tuned LVLM to enable alignment of various visualization modalities, including chart, text, and sketch; and 4) the visualization interaction module decomposes user queries into retrieval tasks with visualization references and utilizes the LVLM to generate the corresponding trend words to fill the textual responses for users.
2.2 Multi-modal Visualization Alignment
Given that visualizations are inherently stored as images, we opt to leverage emerging vision-language models to embed visualization images and facilitate multi-modal user interactions. In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA[36, 37, 19], chart retrieval[52] and generation[43, 53]. Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen. Furthermore, these models are not specifically designed for effectively identifying data change patterns (see Sect. 6.1), a gap that this work seeks to fill in the context of table reasoning.
To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy. Most existing multi-modal data alignment methods rely on contrastive language-image pre-training (CLIP)[41], which has demonstrated its adaptability and proficiency across various tasks[59, 29, 60], highlighting the effectiveness of joint embedding spaces. However, training an effective CLIP-based alignment model for visualizations is non-trivial. First, this task involves aligning more than two modalities, including text, hand-drawn sketches, and various types of charts, whereas the original CLIP model was designed to support text-image alignment. To encompass more modalities, we adopt joint embedding strategies that leverage charts as naturally aligned supervision and emergent alignments to unify other modalities[16]. This approach allows us to prepare paired data only for chart-text and chart-sketch, rather than other pairings such as sketch-text. Second, the effectiveness of CLIP is largely attributed to its extensive training on vast (natural) image-text datasets, whereas there is a lack of suitable training datasets for the task in this work. Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[23, 36]), chart-chart (e.g.,[34, 28, 55]), chart-sketch (e.g.,[35, 45, 32]) similarity measurement. However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for identifying data change patterns. To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multi-modal visualization alignment model.
2.3 Visualization Storage, Indexing, and Retrieval
To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them. With the recent advancements in Vector Database Management Systems (VDBMS)[39], there is potential to meet our demand. VDBMSs are specifically designed for managing high-dimensional data, such as embeddings, offering sophisticated storage, indexing, query processing, and query optimization capabilities[3]. Their ability to perform efficient similarity searches in high-dimensional spaces makes them particularly suited for tasks involving the retrieval of embeddings based on content or semantic similarity. This supports various applications such as information retrieval, e-commerce, and recommendation platforms[17, 47]. Existing VDBMS can be roughly categorized into two types: native systems and extended systems[39]. The former, such as Chroma[1] and EuclidesDB[2], are tailored for vector data management, while the latter, such as AnalyticDB-V[50], are built on top of an existing data management system.
In this work, we first introduce a visualization pruning method to filter out less informative visualizations, and effectively reduce storage demands while maintaining the quality of stored visualizations. We then integrate Chroma[1] into VisLTR and utilize an efficient graph-based index to store and index the large collection of visualization references. An Approximate k-Nearest Neighbors (ANN) algorithm is incorporated for quick and accurate retrieval of visualizations. The combination of pruning with indexing strategies allows for quick retrieval and handling timely user queries.
3 The Visualization-in-the-Loop Framework
@@
The proposed visualization-in-the-loop framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.

### v2

# VisTR: Visualizations as Representations for Time-series Table Reasoning

@@ Abstract
Table reasoning involves transforming natural language questions into corresponding answers based on the provided data table. Recent research exploits large language models (LLMs) to facilitate table reasoning, which however struggle with pattern recognition and lack support for visual-based pattern exploration.To address these limitations, we propose VisTR, a framework that leverages visualizations as representations to facilitate data pattern recognition and support cross-modal exploration. We describe VisTR as a process consisting of four major modules: 1) visualization alignment that utilizes multimodal LLMs to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multimodal interactions for user-friendly table reasoning. Quantitative evaluation with existing multimodal LLMs demonstrates the effectiveness of the alignment model in cross-modal visualization pairings. We further illustrate the applicability of the proposed framework in various time-series table reasoning and exploration tasks.
@@
Table reasoning, visualization representation, multimodal LLMs
@@
VisTR leverages visualizations as the representations between a table and user intentions. The proposed framework advances existing LLM-based table reasoning methods by supporting data change pattern recognition (Q1), improving pattern recognition for long tables (Q2), and enabling visual-based pattern exploration (Q3).
@@ 1 Introduction
Table reasoning is a crucial natural language processing task that involves automatically extracting and inferring information from tables to answer specific questions[12, 70]. This task supports a wide range of applications such as fact verification[6] and question answering[26], where each question-table pair constitutes an instance for reasoning[69]. Given the superior commonsense and logical reasoning capabilities of large language models (LLMs)[63], utilizing them for table reasoning has emerged as the predominant approach[5, 12]. Existing methods can be broadly categorized into generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as designing prompts[5] to guide iterative reasoning or employing flexible tabular formats[57, 69]; and program-aided table reasoning that utilizes executable query languages (e.g., SQL, SPARQL, or Python) as the bridge between user queries and answers derived from the table[65, 9, 19].
@@
Despite their advancements, there exist several limitations. First, these methods struggle with pattern recognition, which involves the automatic identification and extraction of meaningful patterns within data, such as trends in time-series data[35, 8]. For instance, Q1 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning illustrates a scenario where users seek to examine the price trend of Apple during March. Existing LLM-based methods cannot correctly capture and respond to the data change patterns within the corresponding time period. The challenge becomes more pronounced when dealing with large-size tables, as LLMs are known to encounter ‘drift’ issues, where models lose information in the middle of long contexts[4]. This loss of context can result in errors during the reasoning process, making it difficult to extract patterns accurately, as illustrated in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q2). Last, the methods lack support for visual-based pattern exploration, making it challenging for users to iteratively engage with data tables, particulalry for lay users. Rather than expressing their intentions precisely in natural language, laymen would prefer to use multimodal inputs like hand-drawn sketches or charts[41]. An example is shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3), where a user may not know the term “two-valley” but can easily recognize this visual pattern from the collected bar chart. However, existing LLM-based methods do not support such visual-based exploration, which are increasingly in demand.
@@
To overcome the limitations of existing LLM-based table reasoning methods, we introduce a novel framework that incorporates visualization as representation into multimodal LLMs (MLLMs). Our key idea is to transform data tables into a set of meaningful and insightful visualizations, termed visualization references, which serve as a concise data representation to capture data insights and align with human intentions. This framework offers several advantages that address the shortcomings of existing LLM-based table reasoning methods, effectively enhancing the table exploration process. First, visualization references can help recognize the short- and long-term data change patterns. For instance, when recognizing price trend in the given data table (Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q1)), the corresponding visualization reference could capture the “two-peak” price trend of Apple during March. Second, regardless of the length of data facets, the visualization references can capture and retain the data change patterns in fixed-size images. Adopting fixed-size visualization references as the representations of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning. For Q2 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference. Moreover, leveraging visualizations as a representation offers new modalities for interacting with the tabular data, aligning more intuitively with human intentions[32]. Combined with multimodal input, including text and chart (as shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3)), it becomes easier to explore data tables than directly observing the raw table.
@@
We illustrate the effectiveness of VisTR on time-series data tables, which often feature large table sizes and complex data change patterns. In this context, the proposed framework also presents several challenges to be addressed. First, time-series data often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references. To ensure the visualization references are comprehensive enough to accurately represent a data table, thereby meeting the diverse analytical needs, poses the first challenge (C1). Second, to effectively align textual descriptions, visualization images, and hand-drawn sketches to bridge user intentions with visualization references, meanwhile enabling multimodal interactions presents a challenge (C2). Existing models for vision-language tasks such as chart question answering (CQA)[24, 68] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns. Last, to comprehensively represent a table, a large number of visualization references will be generated. Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge (C3).
@@
We design a prototype system that encompasses four main modules aimed at addressing the challenges C1--C3. First, the visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table. To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1). Second, in the visualization alignment module, we fine-tune a MLLM to connect visualizations of various modalities including charts, text, and sketches (addressing C2). Specifically, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions. We further leverage the CLIP model[48] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment. Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3). Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multimodal interactions. Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisTR with the input table. Specifically, the user input is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned MLLM. Subsequently, these descriptions are filled back into the input table to generate textual responses for the user. Quantitative evaluations demonstrate the effectiveness of the fine-tuned MLLM in aligning different modalities. Additionally, case studies illustrate how the proposed framework addresses the limitations of existing LLM-based table reasoning methods through various time-series table reasoning and visual-based exploration tasks. Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
@@
(1) Framework. We introduce a framework that leverages visualizations as the representations to enhance table reasoning, to facilitate data change pattern recognition and visual-based exploration.
(2) Multimodal Visualization Alignment. We design an effective MLLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches. The model aligns multimodal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
(3) VisTR System. We develop a prototype system that supports the exploration and reasoning of large-scale and multiple time-series tables through intuitive user interactions. Various case studies and quantitative results prove the effectiveness and applicability of VisTR.
@@ 2 Related Work
Table Reasoning. Table reasoning integrated with natural language (NL) interfaces enables lay users to explore tabular data effortlessly, supporting complex tasks such as fact verification[6] and question answering[26]. Existing LLM-based table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[57]. Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[5]. However, due to the complexity of table structure, these methods struggle with large tables and offer limited explainability because of their end-to-end nature. Program-aided table reasoning utilizes executable query languages, such as SQL, to form a structured and accountable bridge between user quires and the derived answers[14, 37, 25]. Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[65], using prompt engineering[30], chain-of-thought[57, 66], and pre-training techniques[56]. However, program-aided approaches encounter intricate limitations due to their reliance on query language. First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[29], overlooking data change patterns. Second, when confronted with large tables, LLMs may encounter drift issues[4] while accessing information in the middle of long contexts. Moreover, in many scenarios, relying solely on textual inputs limits the expression of non-verbal intentions, such as hand-drawn sketches, reducing the flexibility of user interactions.
@@
To overcome these limitations, this work introduces a novel framework that leverages visualizations as a representation between tables and user intentions. Our inspiration is that visualizations can depict intuitive data patterns and illustrate complex underlying processes[18], aiding swift and effective interpretation by users[31]. However, this approach also presents challenges to be tackled. First, tabular data typically encompass numerous features, leading to a plethora of potential visualizations[62, 54] that complicate interactive interactions. Second, there are no direct approaches available to connect free-form user queries with visualizations.
@@
Figure 1: An overview of the proposed framework, which mainly contains four main modules of visualization referencing, visualization pruning, visualization alignment, and visualization interaction.
Multimodal Visualization Alignment. Given that visualizations are inherently stored as images, we opt to leverage emerging MLLMs to embed visualization images and facilitate multimodal user interactions. In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA[42, 43, 24], chart retrieval[60] and generation[50, 61]. Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen. Furthermore, these models are not specifically designed for effectively recognizing data change patterns (see Sect. 6.1), a gap that this work seeks to fill in the context of table reasoning.
To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy. Most existing multimodal data alignment methods rely on contrastive language-image pre-training (CLIP)[48], which has demonstrated its adaptability and proficiency across various tasks[71, 34, 72], highlighting the effectiveness of joint embedding spaces. However, training an effective CLIP-based alignment model for visualizations is non-trivial. First, this task involves aligning more than two modalities, including text, hand-drawn sketches, and various types of charts, whereas the original CLIP model was designed to support text-image alignment. To encompass more modalities, we adopt joint embedding strategies that leverage charts as naturally aligned supervision and emergent alignments to unify other modalities[20]. This approach allows us to prepare paired data only for chart-text and chart-sketch, rather than other pairings such as sketch-text. Second, the effectiveness of CLIP is largely attributed to its extensive training on vast (natural) image-text datasets, whereas there is a lack of suitable training datasets for the task in this work. Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[28, 42]), chart-chart (e.g.,[40, 33, 64]), chart-sketch (e.g.,[41, 52, 38]) similarity measurement. However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for recognizing data change patterns. To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multimodal visualization alignment model.
Visualization Storage, Indexing, and Retrieval. To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them. With the recent advancements in Vector Database Management Systems (VDBMS)[46], there is potential to meet our demand. VDBMSs are specifically designed for managing high-dimensional data, such as embeddings, offering sophisticated storage, indexing, query processing, and query optimization capabilities[3]. Their ability to perform efficient similarity searches in high-dimensional spaces makes them particularly suited for tasks involving the retrieval of embeddings based on content or semantic similarity. This supports various applications such as information retrieval, e-commerce, and recommendation platforms[22, 55]. Existing VDBMS can be roughly categorized into two types: native systems and extended systems[46]. The former, such as Chroma[1] and EuclidesDB[2], are tailored for vector data management, while the latter, such as AnalyticDB-V[58], are built on top of an existing data management system.
In this work, we first introduce a visualization pruning method to filter out less informative visualizations, and effectively reduce storage demands while maintaining the quality of stored visualizations. We then integrate Chroma[1] into VisTR and utilize an efficient graph-based index to store and index the large collection of visualization references. An Approximate k-Nearest Neighbors (ANN) algorithm is incorporated to retrieve visualizations. The combination of pruning with indexing allows for quick retrieval and handling timely user queries.
3 The Visualization Enhanced Framework
The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.